OpenStudy (anonymous):

a-9 a^2-3 2- ---- = ------ a+4 a +4 solve

OpenStudy (anonymous):

common numerator on a right site: [2(a+4) - (a-9)]/(a+4) = (a^2-3)/(a+4) do you see that denominator is the same on both sides.. which is means that

OpenStudy (anonymous):

numerators should be the same. so we can write that following: 2a+8-a+9=a^2-3 a^2-a-20=0 solve this quadratic equation for a

OpenStudy (anonymous):

(x+4)(x-5)=0 x=-4 and x=5 the root a=-4 can not be the answer because you can delete by 0, so there is only one answer: x=5